OpenStudy (anonymous):

Percentage Yield problem please?? Consider the unbalanced equation for the decomposition of silver nitrate: AgNO3(s) --->Ag(s)+NO2(g)+O2(g) 1. What is the percent yield of Ag for the reaction?

OpenStudy (anonymous):

This is continue step for your previous post: Theoretical mole of AgNO3: 54/ 170 = .318 mol Theoretical mass of Ag: .318 * 108 = 34.31g Percent yield: 31.05 * 100 / 34.31 = 90.51%